# Fast nonparametric feature selection with error control using integrated path stability selection

### v1

@@ Abstract.
Feature selection can greatly improve performance and interpretability in machine learning problems. However, existing nonparametric feature selection methods either lack theoretical error control or fail to accurately control errors in practice. Many methods are also slow, especially in high dimensions. In this paper, we introduce a general feature selection method that applies integrated path stability selection to thresholding to control false positives and the false discovery rate. The method also estimates $q$-values, which are better suited to high-dimensional data than $p$-values. We focus on two special cases of the general method based on gradient boosting (IPSSGB) and random forests (IPSSRF). Extensive simulations with RNA sequencing data show that IPSSGB and IPSSRF have better error control, detect more true positives, and are faster than existing methods. We also use both methods to detect microRNAs and genes related to ovarian cancer, finding that they make better predictions with fewer features than other methods.
@@ 1.  Introduction
Identifying the important features in a dataset can greatly improve performance and interpretability in machine learning problems (Theng and Bhoyar, 2024). For example, in genomics, often only a small fraction of genes (features) are related to a disease of interest (response). By identifying these genes, scientists can save time and resources while gaining insights that would be difficult to uncover otherwise (Theng and Bhoyar, 2024). A common form of feature selection is thresholding: Given a function $\Phi$ that assigns a degree of importance $\Phi(j)$ to each feature $j$ in a dataset, one seeks a value $\lambda$ such that the selected set $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$ contains important features and its complement, $\hat{S}^{c}_{\lambda}=\{j:\Phi(j)<\lambda\}$, does not. Choosing $\lambda$ can be challenging, and many methods for doing so come with few theoretical guarantees (Huynh-Thu et al., 2012).
@@
Stability selection is a popular technique for improving the performance of feature selection algorithms (Meinshausen and Bühlmann, 2010). However, its implementation relies on relatively weak upper bounds on the expected number of false positives, E(FP), that is, on the expected number of unimportant features that are selected. Recently, Melikechi and Miller (2024) proved that much stronger bounds hold for integrated path stability selection (IPSS), which consequently identifies more true positives in practice than stability selection. Until now, IPSS has only been applied to parametric linear models, and its generalization to more complex models has not been explored.
@@
In this work, we extend IPSS to nonlinear, nonparametric models by applying it to thresholding. The result is a feature selection method with tight upper bounds on E(FP) that are characterized by novel quantities called efp scores. In addition to controlling E(FP), efp scores also approximately control the false discovery rate (FDR) and estimate the $q$-values for each feature, which are more reliable and robust than $p$-values in genomics and other high-dimensional settings (Storey and Tibshirani, 2003). Lastly, by integrating over threshold values, IPSS for thresholding replaces the problem of specifying a single threshold $\lambda$ with the simpler task of specifying either a target E(FP) or a target FDR.
@@
We develop two specific instances of the general method: IPSS for gradient boosting (IPSSGB) and IPSS for random forests (IPSSRF). In simulations, we find that IPSSGB has better false positive control and identifies more true positives than nine other feature selection methods, and that both IPSSGB and IPSSRF outperform IPSS for linear models when the linearity assumptions of the latter are violated. More generally, IPSSGB and IPSSRF control E(FP) and FDR more accurately than other stability selection and knockoff-based methods with theoretical error control. In Section 5, we find that IPSSGB and IPSSRF effectively identify microRNAs and genes related to ovarian cancer, achieving better predictive performance than other feature selection methods while using fewer features. Both also run faster than leading methods.
Organization. In Section 2, we introduce efp scores, IPSS for thresholding in general, and IPSS for gradient boosting and random forests. In Section 3, we describe several leading feature selection methods, which are compared to IPSSGB and IPSSRF in a simulation study in Section 4. In Section 5, we analyze ovarian cancer data. We conclude in Section 6 with a discussion.
@@ 2.  Methods
We introduce efp scores (Section 2.1), review IPSS (Section 2.2), introduce IPSS for thresholding (Section 2.3), and apply IPSS to gradient boosting and random forests (Section 2.4).
@@
2.1.  efp scores.
Suppose $S\subseteq\{1,\ldots,p\}$ is an unknown subset of true important features that we wish to estimate using $Z_{1:n}$. An efp (expected false positive) score is a function $\mathtt{efp}_{Z_{1:n}}:\{1,\dots,p\}\to[0,\infty)$ that depends on $Z_{1:n}$ and satisfies the following:
@@
E(FP) and efp scores are related to two other quantities of significant interest: the false discovery rate (FDR) and $q$-values. Informally, the false discovery rate is the expected ratio between the number of false positives and the total number of features selected, $\mathrm{FDR}=\mathrm{E(FP/(TP+FP))}$, and the $q$-value of feature $j$ is the smallest FDR when $j$ is selected (Storey, 2003). When $p$ is large, as is often the case in genomics, Storey (2003) showed that
@@
where $\mathrm{pFDR}(t)=\mathrm{E}(\mathrm{FP}(t)/\lvert\hat{S}(t)\rvert\mid\lvert\hat{S}(t)\rvert>0)$ is the positive false discovery rate, and the inequality holds by the definition of an efp score. It follows that the $q$-value of feature $j$ satisfies
@@
2.2.  Integrated path stability selection.
In Section 2.1, we showed that efp scores can be used to control E(FP), the FDR, and $q$-values. In this section, we review integrated path stability selection (IPSS), which constructs efp scores from arbitrary feature selection algorithms. The advantage of IPSS over other stability selection methods is that its efp scores have the tightest E(FP) bounds (Melikechi and Miller, 2024).
Let $S$ be the unknown subset of important features as before, and let $\hat{S}_{\lambda}\subseteq\{1,\ldots,p\}$ be an estimator of $S$ that depends on the data and a parameter $\lambda>0$, with larger values of $\lambda$ corresponding to fewer features being selected. Importantly, $\hat{S}_{\lambda}$ and $\hat{S}(t)$ are distinct estimators of $S$: The former is a baseline selection algorithm whose parameter $\lambda$ will always appear as a subscript, and the latter is an estimator based on efp scores whose parameter $t$ will always appear as a functional argument.
Stability selection uses subsampling to construct $\hat{S}(t)$ from $\hat{S}_{\lambda}$ (Meinshausen and Bühlmann, 2010; Shah and Samworth, 2013). Specifically, rather than estimate $S$ using all of the data at once, one randomly draws disjoint subsets $A_{1},A_{2}\subseteq\{1,\dots,n\}$ of size $\lfloor n/2\rfloor$ and evaluates $\hat{S}_{\lambda}(Z_{A_{1}})$ and $\hat{S}_{\lambda}(Z_{A_{2}})$ at all $\lambda$ in some interval $\Lambda\subseteq(0,\infty)$, where $Z_{A}=(Z_{i}:i\in A)$. After $B$ subsampling iterations, the estimated selection probability $\hat{\pi}_{j}(\lambda)$ is the proportion of times feature $j$ is selected over all $2B$ subsets, $\hat{\pi}_{j}(\lambda)=\frac{1}{2B}\sum_{b=1}^{2B}\operatorname{\mathds{1}}(j\in\hat{S}_{\lambda}(Z_{A_{b}}))$.
@@
is a valid efp score, where $\mathcal{I}(\Lambda)$ is a parameter that is described below and in Section S1. While several choices of $f$ are possible in Equation 2.3, we always use $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ because the resulting efp score bounds $\mathrm{E}(\mathrm{FP}(t))\leq t$ are the tightest of any existing version of stability selection (Melikechi and Miller, 2024).
@@
2.3.  IPSS for thresholding.
An importance function is a (possibly random) map $\Phi_{Z_{1:n}}:\{1,\dots,p\}\to\mathbb{R}$ that uses the data to assign an importance score $\Phi_{Z_{1:n}}(j)$ to each feature. The possible randomness, which is additional to the randomness in $Z_{1:n}$, can come from, for example, random subsampling in tree-based algorithms. Suppressing $Z_{1:n}$ from the notation for now, assume that $\Phi(j)<\Phi(k)$ means $j$ is less important than $k$ according to $\Phi$. For example, in linear regression where $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}=\beta^{\mathtt{T}}X_{i}+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, a viable importance function is the magnitude of the estimated regression coefficient, $\Phi(j)=\lvert\hat{\beta}_{j}\rvert$. Another class of importance functions—the focus in this work—come from tree ensemble methods such as gradient boosted trees and random forests, where again we are in the supervised setting $Z_{i}=(X_{i},Y_{i})$ (Friedman, 2001; Breiman, 2001). Given a collection of binary decision trees, $\mathcal{T}$, these importance functions take the form
@@
where the outer sum is over all trees $T\in\mathcal{T}$, the inner sum is over all nodes $v\in T$, $j_{v}$ is the feature used to split node $v$, and $\varphi(v)$ measures the impurity of $v$. For regression, we use the squared error loss, $\varphi(v)=\sum_{i\in v}(Y_{i}-\overline{Y}_{v})^{2}/\lvert v\rvert$, where $v\subseteq\{1,\dots,n\}$ is identified with the subset of samples in node $v$, and $\overline{Y}_{v}=\sum_{i\in v}Y_{i}/\lvert v\rvert$ is the empirical mean of the responses in $v$. For binary classification, we use the Gini index, $\varphi(v)=2p_{0}(v)p_{1}(v)$ where, for $a\in\{0,1\}$, $p_{a}(v)=\sum_{i\in v}\operatorname{\mathds{1}}(Y_{i}=a)/\lvert v\rvert$ is the proportion of responses in $v$ that equal $a$. The change in impurity

### v2

# Nonparametric IPSS Fast, flexible feature selection with false discovery control

@@ Abstract.
Feature selection is a critical task in machine learning and statistics. However, existing feature selection methods either (i) rely on parametric methods such as linear or generalized linear models, (ii) lack theoretical false discovery control, or (iii) identify few true positives. Here, we introduce a general feature selection method with finite-sample false discovery control based on applying integrated path stability selection (IPSS) to arbitrary feature importance scores. The method is nonparametric whenever the importance scores are nonparametric, and it estimates $q$-values, which are better suited to high-dimensional data than $p$-values. We focus on two special cases using importance scores from gradient boosting (IPSSGB) and random forests (IPSSRF). Extensive nonlinear simulations with RNA sequencing data show that both methods accurately control the false discovery rate and detect more true positives than existing methods. Both methods are also efficient, running in under 20 seconds when there are 500 samples and 5000 features. We apply IPSSGB and IPSSRF to detect microRNAs and genes related to cancer, finding that they yield better predictions with fewer features than existing approaches.
@@ 1.  Introduction
Identifying the important features in a dataset can greatly improve performance and interpretability in machine learning and statistical problems (Theng and Bhoyar, 2024). For example, in genomics, often only a small fraction of genes (features) are related to a disease of interest (response). By identifying these genes, scientists can save time and resources while gaining insights that would be difficult to uncover otherwise (Theng and Bhoyar, 2024).
The goal of feature selection is to maximize the number of important features selected (true positives)—informally referred to here as power—while minimizing the number of unimportant features selected (false positives). In simulations, we find that popular methods without theoretical false discovery control, such as recursive feature elimination and Boruta (Kursa et al., 2010), often select many false positives (Section 3). Meanwhile, popular methods with false discovery control, namely stability selection (Meinshausen and Bühlmann, 2010; Shah and Samworth, 2013) and model-X knockoffs (Candes et al., 2018), have low power in simulations and select few features in practice (Sections 3 and 4).
@@
One reason for stability selection’s low power is its relatively weak theoretical upper bounds on the expected number of false positives, E(FP). Recently, Melikechi and Miller (2024) proved that much stronger bounds hold for integrated path stability selection (IPSS), which consequently identifies more true positives. Until now, IPSS has only been applied to generalized linear models, limiting its applicability since parametric assumptions are often violated or difficult to verify in practice.
@@
Thus, there is a need for nonparametric feature selection methods with theoretical false discovery control and greater power. In this work, we address this need by applying IPSS to arbitrary feature importance scores. The result is a general feature selection method with tight upper bounds on E(FP) that are characterized by novel quantities called efp scores. In addition to controlling E(FP), efp scores approximately control the false discovery rate (FDR) and estimate $q$-values for each feature, which are more reliable than $p$-values in genomics and other high-dimensional settings (Storey and Tibshirani, 2003).
Our proposed method is nonparametric whenever the feature importance scores come from nonparametric models. We develop two specific instances of this: IPSS for gradient boosting (IPSSGB) and IPSS for random forests (IPSSRF). Like knockoffs, neither IPSSGB nor IPSSRF assume a specific functional relationship between the response and the features. Unlike knockoffs, neither method requires knowledge of the joint distribution of the features.
@@
In simulations, we find that IPSSGB and IPSSRF provide a better balance of FDR control and power than 12 other methods, and that IPSSGB performs best overall. In particular, both methods significantly outperform parametric versions of IPSS when the parametric assumptions are violated. In Section 4, we find that IPSSGB and IPSSRF successfully identify microRNAs and genes related to ovarian cancer and glioma, achieving better predictive performance than other feature selection methods while using fewer features. Both are also computationally efficient.
Finally, another important aspect of feature selection is stability—the consistent selection of features across similar settings. Stability improves reproducibility, which is critical in many applications (Li et al., 2022). As their names suggest, stability selection and IPSS are designed to produce stable results by repeatedly applying a baseline feature selection algorithm to random subsamples of the data (Section 2.2). Nogueira et al. (2018) show that stability selection can produce significantly more stable results than its baseline algorithm. In this work, we focus on false discovery control and power; further study of the stability of IPSS, stability selection, and other stability-inspired methods like StabML-RFE (Li et al., 2022) is left to future work.
Organization. In Section 2, we introduce efp scores, review IPSS, and present its extension to feature importance scores. In Section 3, we present our simulation studies, and in Section 4, we analyze ovarian cancer and glioma data. We conclude in Section 5 with a discussion.
@@ 2.  Methods
In Section 2.1, we introduce efp scores. These quantities, assigned to each feature in the dataset, are used to perform feature selection with E(FP) control. In Section 2.2, we introduce IPSS, which is a general approach for constructing efp scores. In Section 2.3, we describe how IPSS can be applied to any feature importance score, focusing in particular on importance scores from tree-based methods such as gradient boosting and random forests.
@@
2.1  efp scores.
Suppose $S\subseteq\{1,\ldots,p\}$ is an unknown subset of important features that we wish to estimate using $Z_{1:n}$. An efp (expected false positive) score is a function $\mathtt{efp}_{Z_{1:n}}:\{1,\dots,p\}\to[0,\infty)$ that depends on $Z_{1:n}$ and satisfies the following:
@@
E(FP) and efp scores are related to two other quantities of significant interest: the false discovery rate (FDR) and $q$-values. Informally, the false discovery rate is the expected ratio between the number of false positives and the total number of features selected, $\mathrm{FDR}=\mathrm{E(FP/(TP+FP))}$, and the $q$-value of feature $j$ is the smallest FDR when $j$ is selected (Storey, 2003). When $p$ is large, as is often the case in genomics, we have
@@
where $\mathrm{pFDR}(t)=\mathrm{E}(\mathrm{FP}(t)/\lvert\hat{S}(t)\rvert\mid\lvert\hat{S}(t)\rvert>0)$ is the positive false discovery rate, the two approximations are from Storey (2003), and the inequality holds by the definition of an efp score. It follows that the $q$-value of feature $j$ satisfies
@@
2.2  Integrated path stability selection.
Integrated path stability selection (IPSS) constructs efp scores by applying baseline feature selection algorithms to random subsamples of the data. Specifically, let $S$ be an unknown subset of true features as before, and let $\hat{S}_{\lambda}\subseteq\{1,\ldots,p\}$ be an estimator of $S$ that depends on the data and a parameter $\lambda>0$. Note that $\hat{S}_{\lambda}$ and $\hat{S}(t)$ are distinct estimators of $S$: The former is a baseline algorithm whose parameter $\lambda$ appears as a subscript.
The IPSS subsampling procedure, which is also used in stability selection, consists of $B$ subsampling iterations. Each iteration consists of randomly drawing disjoint subsets $A_{1},A_{2}\subseteq\{1,\dots,n\}$ of size $\lfloor n/2\rfloor$ and evaluating $\hat{S}_{\lambda}(Z_{A_{1}})$ and $\hat{S}_{\lambda}(Z_{A_{2}})$ at all $\lambda$ in some interval $\Lambda\subseteq(0,\infty)$, where $Z_{A}=(Z_{i}:i\in A)$ (the choice of $\lfloor n/2\rfloor$ samples is needed for existing stability selection theorems—not just IPSS—to hold). After $B$ iterations, the estimated selection probability $\hat{\pi}_{j}(\lambda)=\frac{1}{2B}\sum_{b=1}^{2B}\operatorname{\mathds{1}}(j\in\hat{S}_{\lambda}(Z_{A_{b}}))$ of feature $j$ is the proportion of times $j$ is selected over all $2B$ subsets. Large values of $\hat{\pi}_{j}(\lambda)$ correspond to $j$ being selected by $\hat{S}_{\lambda}$ on many of the random subsamples, suggesting that $j$ is important.
@@
is a valid efp score, that is, $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$ satisfies $\mathrm{E}\lvert\hat{S}(t)\cap S^{c}\rvert\leq t$. An explicit form of $\mathcal{I}(\Lambda)$, which comes from the theory of IPSS, is provided in Section S1.
Several choices of $f$ in Equation 2.3 yield provably valid efp scores (see Section S1 as well as Theorems 4.1, 4.2, S3.2, S3.3 in Melikechi and Miller (2024) for detailed theoretical results about IPSS with different functions). We always use $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ because the resulting bounds $\mathrm{E}(\mathrm{FP}(t))\leq t$ are the tightest of any existing version of stability selection (Melikechi and Miller, 2024). In addition to its theoretical justification, the empirical results in Figures S18 and S19 show that this choice of $f$ produces the best results among several functions for which valid efp scores are available. Further details about IPSS, including descriptions of $\Lambda$, $\mu$, and a derivation of Equation 2.3, are provided in Section S1.
2.3  IPSS for feature importance scores.
Until now, IPSS has only been applied to parametric baseline estimators $\hat{S}_{\lambda}$. A canonical example is $\ell^{1}$-regularized estimators $\hat{S}_{\lambda}=\{j:\hat{\beta}_{j}(\lambda)\neq 0\}$, where
$\hat{\beta}(\lambda)$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\sum_{i=1}^{n}\mathcal{L}(Y_{i},\beta^{\mathtt{T}}X_{i})+\lambda\sum_{j=1}^{p}\lvert\beta_{j}\rvert.$ (2.4)
Here, $\lambda>0$ controls the strength of the $\ell^{1}$-penalty and $\mathcal{L}$ is a log-likelihood that assumes a specific relationship between the features and response. By extending IPSS to feature importance scores, we no longer require such restrictive assumptions.
An importance function is a (possibly random) map $\Phi_{Z_{1:n}}:\{1,\dots,p\}\to\mathbb{R}$ that uses the data to assign an importance score $\Phi_{Z_{1:n}}(j)$ to each feature. The possible randomness, which is additional to the randomness in $Z_{1:n}$, can come from, for example, random subsampling in tree-based algorithms. Suppressing $Z_{1:n}$ from the notation for now, assume $\Phi(j)<\Phi(k)$ means that $j$ is less important than $k$ according to $\Phi$. For example, in linear regression where $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}=\beta^{\mathtt{T}}X_{i}+\epsilon_{i}$, a viable importance function is the magnitude of the estimated regression coefficient, $\Phi(j)=\lvert\hat{\beta}_{j}\rvert$.
Associated to any importance function $\Phi$ is a baseline feature selection estimator $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$, obtained by simply thresholding the importance scores. That is, given $\Phi$ and a threshold $\lambda$, $\hat{S}_{\lambda}$ selects the features whose importance scores are at least $\lambda$. With $\hat{S}_{\lambda}$ defined, we have all that is needed to implement IPSS. Furthermore, all of the theoretical results from Melikechi and Miller (2024) apply without modification.
2.3.1  Implementation.
Algorithm S1 outlines IPSS for feature importances. The main steps are as follows: First, features are preselected according to the procedure described in Section S2.1. This is a common preliminary step in many feature selection algorithms. Next, importance scores are evaluated for the preselected features using random, disjoint halves of the data. This process is repeated $B$ times, yielding $2B$ importance scores for each feature. We then construct a grid of $\lambda$ values. The largest, $\lambda_{\mathrm{max}}$, is the maximum importance score across all features and all $2B$ sets of scores (hence, $\hat{S}_{\lambda_{\mathrm{max}}}=\varnothing$). Next, starting from $\lambda_{\mathrm{min}}=\lambda_{\mathrm{max}}$, decrease $\lambda_{\mathrm{min}}$ one step at a time, usually on a log scale, iteratively updating the integral $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ at each step in the form of a Riemann sum approximation until $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ surpasses a cutoff $C$. Once $C$ is surpassed, the $\mathtt{while}$ loop stops and the feature-specific selection probabilities and integrals are evaluated. The algorithm outputs efp scores for each feature, which are used to control E(FP), FDR, and compute $q$-values, as described in Section 2.1.
@@
2.3.2  IPSSGB and IPSSRF.
Any importance function can be combined with IPSS, providing many possible directions for future work. In this paper, we focus on importance functions from tree ensemble methods because they are nonparametric, computationally efficient, and produce state-of-the-art results on tabular data (Shwartz-Ziv and Armon, 2022). These importance functions are defined as follows; for additional details, see Louppe et al. (2013) or Biau and Scornet (2016).
Given a collection of binary decision trees, $\mathcal{T}$, define
@@
where the outer sum is over all trees $T\in\mathcal{T}$, the inner sum is over all nodes $v\in T$, $j_{v}$ is the feature used to split node $v$, and $\varphi(v)$ measures the impurity of $v$. The change in impurity